# Insights from leptohadronic modelling of the brightest blazar flare

### v1

@@ Abstract
The blazar 3C 454.3 experienced a major flare in November 2010 making it the brightest $\gamma$-ray source in the sky of the Fermi-LAT. We obtain seven daily consecutive spectral-energy distributions (SEDs) of the flare in the infra-red, optical, ultra-violet, X-ray and $\gamma$-ray bands with publicly available data. We simulate the physical conditions in the blazar and show that the observed SEDs are well reproduced in the framework of a “standing feature” where the position of the emitting region is almost stationary, located beyond the outer radius of the broad-line region and into which fresh blobs of relativistically moving magnetized plasma are continuously injected. Meanwhile, a model with a single “moving blob” does not describe the data well. We obtain a robust upper limit to the amount of high-energy protons in the jet of 3C 454.3 from the electromagnetic SED. We construct a neutrino light curve of 3C 454.3 and estimate the expected neutrino yield at energies $\geq 100$ TeV for 3C 454.3 to be up to $6\times 10^{-3}$ $\nu_{\mu}$ per year. Finally, we extrapolate our model findings to the light curves of all Fermi-LAT flat-spectrum radio quasars. We find that next-generation neutrino telescopes are expected to detect approximately one multimessenger ($\gamma+\nu_{\mu}$) flare per year from bright blazars with neutrino peak energy in the hundreds TeV – hundreds PeV energy range and show that the electromagnetic flare peak can precede the neutrino arrival by months to years.
@@
pubyear: 2025pagerange: Insights from leptohadronic modelling of the brightest blazar flareE
@@ 1 Introduction
Blazars – bright active galactic nuclei (AGNs) with their relativistic jets pointing at small angles towards Earth – dominate the $\gamma$-ray sky visible by the Fermi Large Area Telescope (Fermi-LAT) (Atwood et al., 2009) in the region far from the galactic plane (Abdollahi et al., 2020). Since the evidence of the association of the IceCube neutrino event IceCube-170922A with the electromagnetic flare of blazar TXS 0506+056 in 2017 (IceCube Collaboration et al., 2018), blazars became objects of intensive theoretical studies as potential sources of the astrophysical neutrinos (e.g., Keivani et al., 2018; Ansoldi et al., 2018; Murase et al., 2018; Sahakyan, 2018; Gao et al., 2019; Oikonomou et al., 2019; Cerruti et al., 2019; Reimer et al., 2019; Rodrigues et al., 2019; Liu et al., 2019; Petropoulou et al., 2020; Viale et al., 2024; Robinson & Böttcher, 2024). Multiple searches for correlations between the IceCube neutrinos and different electromagnetic counterparts among blazars resulted in the conclusion that blazars contribute to the total IceCube astrophysical neutrino flux at the level of around $\lesssim 1-30$ (for a review, see Troitsky, 2021, table 4 and references therein; Abbasi et al., 2023b; Troitsky, 2024).
@@
FSRQs are sources with high flaring activity (Ackermann et al., 2011; Abdollahi et al., 2017; Yoshida et al., 2023) with enhanced observable electromagnetic intensity due to the increase of the luminosity of charged particles in the source and/or the apparent Doppler factor which leads to the amplified neutrino production. Depending on the details of a particular model flares may dominantly contribute to the diffuse neutrino flux (Murase et al., 2018). Besides, the atmospheric background is limited by the time window of a flare.
@@
Using publicly-available data from SMARTS, Swift-UVOT, Swift-XRT, and Fermi-LAT, we obtain seven multi-wavelength quasi-simultaneous one-day averaged SEDs observed during the flare period. We describe the multi-wavelength data analysis in Sect. 2. We model (Sect. 3) the observable flare SEDs with the publicly-available AM3 program (Klinger et al., 2023, 2024) within three different frameworks: 1) assuming a standard leptonic model where the observed electromagnetic emission comes from relativistic electrons losing energy via synchrotron and inverse-Compton (IC) radiation in compact blobs passing through a quasi-stationary feature in the jet into which relativistic plasma is injected with large bulk Lorentz factor (“standing-feature” model, Sects. 3.1, 4.1); 2) utilizing the AM3’s capabilities of the time-dependent modelling, we attempt to model the whole flare in framework of a “moving-blob” model where the whole radiation comes from a single blob co-moving with the bulk motion of the relativistic jet plasma (“moving-blob” model, Sects. 3.2, 4.2); 3) assuming that in addition to electrons, protons might be co-accelerated in the jet with the same efficiency, we calculate the maximum-allowed by the data proton energy density $u^{\prime\mathrm{UL}}_{p}$ within a leptohadronic “standing-feature” model (Sec. 3.3, 4.3), providing our method to estimate the expected IceCube neutrino yield in Sect. 3.4 with corresponding results in Sect. 4.4. Our method to extrapolate the results for 3C 454.3 on the whole population of Fermi-LAT FSRQs is presented in Sect. 3.5 and the results of this are presented in Sect. 4.5. We discuss our findings in Sect. 5. In Sect. 6, we conclude that it is not possible to describe the seven days of the flare with a single “moving-blob” model while a “standing-feature” model does it successfully. Moreover, we find that multi-wavelength flares of Fermi-LAT FSRQs similar to the studied 3C 454.3 flare contribute to the neutrino flux at energies $\geq 100$ TeV at the level of $\sim 0.5$% under the assumption that their emission can be described within a single-zone “standing-feature” leptohadronic model with the scaling between the observed integral neutrino and $\gamma$-ray fluxes inferred from the 3C 454.3 modelling for MJD 55517–55524.
@@ 2.1 Fermi-LAT – $\gamma$ rays
For our object of interest, 3C 454.3, we select the Fermi-LAT data in the energy range $100$ MeV $\leq E\leq$ $1$ TeV and in time intervals corresponding to one day within the period MJD 55517–55524. The analysis is performed with the standard fermipy (Wood et al., 2017) package for the binned likelihood analysis. During the Fermi-LAT data analysis of the $15^{\circ}$ ROI (region of interest) centred at the position of 3C 454.3 we apply a cut on the zenith angle value $\theta_{z}\leq 90^{\circ}$ in order to remove contamination from the Earth’s limb. We use the P8R3_SOURCE_V3 instrument response function (evclass $=128$) with detected $\gamma$ rays both in FRONT and BACK of the Fermi-LAT (evtype $=3$) [^1] https://fermi.gsfc.nasa.gov/ssc/data/analysis/documentation/Cicerone/Cicerone_LAT_IRFs/IRF_overview.html. We divide the data into six bins per decade in energy, and the angular binning is $0.1^{\circ}$ per pixel. To describe the observed $\gamma$-ray emission we construct a model containing all sources from the 4th Fermi-LAT source catalogue 4FGL (Abdollahi et al., 2020) located within $15^{\circ}$ from the centre of the ROI2223C 454.3 is present itself in the 4FGL with the name 4FGL J2253.9+1609., the model of the diffuse galactic background gll_iem_v07, and the model of the isotropic $\gamma$-ray background iso_P8R3_SOURCE_V3_v1333https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html. The normalizations of the diffuse backgrounds, 3C 454.3 and other sources within $7.5^{\circ}$ from the centre of the ROI are left free during the initial fitting of the ROI with all spectral shapes fixed according to the catalogue; both normalizations and spectral shapes of other sources outside $7.5^{\circ}$ from the ROI centre are fixed to their catalogue values. After the initial fit, the spectral shape of 3C 454.3 is freed and the fit is repeated. The $\gamma$-ray SED of 3C 454.3 is then obtained using the fermipy.GTAnalysis.sed method with the default parameter values [^4] https://fermipy.readthedocs.io/en/latest/advanced/sed.html#id3.
@@ 2.2 Swift-XRT – X rays
To better constrain the column density along the line of sight (LoS), we first obtain the SED averaged over the whole seven-day period MJD 55517–55524. The observed spectrum is modelled as a broken power law accounting for the photoelectric absorption along the LoS. The solar abundance is adopted from Wilms et al. (2000) and the photoionization cross-section table $\sigma_{\mathrm{photoel}}(E_{i})$ is from Verner et al. (1996). The initial guess for the hydrogen column density $\kappa$ is set to $\kappa=6.81\times 10^{20}$ cm-2 according to the average value calculated with the use of the web tool [^6] https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3nh/w3nh.pl? within the $0.1^{\circ}$-cone with the centre at the position of 3C 454.3 on H I maps obtained by HI4PI Collaboration et al. (2016), although the column density is left free along with the normalization, break energy, and the two spectral indices of the broken-power-law X-ray spectrum. Using fit, plot_unfold, and write_plot ISIS functions we extract both the observed $\mathrm{SED}_{o}(E_{i})$ and the deabsorbed $\mathrm{SED}_{d}(E_{i})$ from the best-fit model and take into account the uncertainty related to the deabsorption correction in the following manner:
$\sigma_{\mathrm{photoel}}(E_{i})=-\frac{\ln k_{a}(E_{i})}{\kappa},$ (3)
where $\varsigma_{d}(E_{i})$ is the final uncertainty of the deabsorbed SED after taking into consideration the uncertainty of the hydrogen column density $\varsigma_{\kappa}$, and $\varsigma_{o}(E_{i})$ is the statistical uncertainty of the observed $\mathrm{SED}_{o}(E_{i})$ in the $i$th energy bin $E_{i}$. $\varsigma_{\kappa}$ is estimated using conf_loop ISIS function calculating 68% confidence interval for each of the free parameters after having obtained the best-fit with the absorbed broken-power-law model. After obtaining $\kappa=7.7^{\,+1.8}_{-2.4}\times 10^{20}$ cm-2 for MJD 55517–55524, we perform the same fit as described in this paragraph for each of the seven one-day-averaged consecutive SEDs allowing $\kappa$ to vary only within the 68% confidence interval obtained above.
@@ 2.3.1 Swift-UVOT
The Swift-UVOT SEDs are obtained with swift-uvot-analysis-tools [^7] https://github.com/KarlenS/swift-uvot-analysis-tools program providing the values of the observed SED averaged over a given day within the time interval MJD 55517–55524.
@@ 2.3.3 Deredenning
The deredenning for IR, optical and UV data was performed using extinction program (Barbary, 2016) assuming the extinction law of Fitzpatrick (1999), with the mean value of $E(B-V)$ extracted from IRSA (2022) using astroquery.ipac.irsa.irsa_dust module (Ginsburg et al., 2024, 2019) with the assumed value of $R_{V}=3.1$.

### v2

@@ Abstract
The blazar 3C 454.3 experienced a major flare in November 2010, making it the brightest $\gamma$-ray source in the sky of the Fermi Large Area Telescope (LAT). We obtain seven daily consecutive spectral-energy distributions (SEDs) of the flare in the infrared, optical, ultraviolet, X-ray and $\gamma$-ray bands with publicly available data. We simulate the physical conditions in the blazar and show that the observed SEDs are well reproduced in the framework of a “standing feature” where the position of the emitting region is almost stationary, located beyond the outer radius of the broad-line region and into which fresh blobs of relativistically moving magnetised plasma are continuously injected. Meanwhile, a model with a single “moving blob” does not describe the data well. We obtain a robust upper limit to the amount of high-energy protons in the jet of 3C 454.3 from the electromagnetic SED. We construct a neutrino light curve of 3C 454.3 and estimate the expected neutrino yield at energies $\geq 100$ TeV for 3C 454.3 to be up to $6\times 10^{-3}$ $\nu_{\mu}$ per year. Finally, we extrapolate our model findings to the light curves of all Fermi-LAT flat-spectrum radio quasars. We find that next-generation neutrino telescopes are expected to detect approximately one multimessenger ($\gamma+\nu_{\mu}$) flare per year from bright blazars with neutrino peak energy in the hundreds TeV – hundreds PeV energy range and show that the electromagnetic flare peak can precede the neutrino arrival by months to years.
@@
pubyear: 2025pagerange: Insights from leptohadronic modelling of the brightest blazar flareF
@@ 1 Introduction
Blazars – bright active galactic nuclei (AGNs) with their relativistic jets pointing at small angles towards Earth – dominate the $\gamma$-ray sky visible by the Fermi Large Area Telescope (Fermi-LAT) (Atwood et al., 2009) in the region far from the galactic plane (Abdollahi et al., 2020). Since the evidence of the association of the IceCube neutrino event IceCube-170922A with the electromagnetic flare of blazar TXS 0506+056 in 2017 (IceCube Collaboration et al., 2018), blazars became objects of intensive theoretical studies as potential sources of the astrophysical neutrinos (e.g., Keivani et al., 2018; Ansoldi et al., 2018; Murase et al., 2018; Sahakyan, 2018; Gao et al., 2019; Oikonomou et al., 2019; Cerruti et al., 2019; Reimer et al., 2019; Rodrigues et al., 2019; Liu et al., 2019; Jiménez-Fernández & van Eerten, 2021; Petropoulou et al., 2020; Das et al., 2021, 2022; Kalashev et al., 2023; Viale et al., 2024; Robinson & Böttcher, 2024). Multiple searches for correlations between the IceCube neutrinos and different electromagnetic counterparts among blazars resulted in the conclusion that blazars contribute to the total IceCube astrophysical neutrino flux at the level of around $\lesssim 1-30$ % (for a review, see Troitsky, 2021, table 4 and references therein; Abbasi et al., 2023b; Troitsky, 2024).
@@
FSRQs are sources with high flaring activity (Ackermann et al., 2011; Abdollahi et al., 2017; Yoshida et al., 2023) with enhanced observable electromagnetic intensity due to the increase of the luminosity of charged particles in the source and/or the apparent Doppler factor, which leads to the amplified neutrino production. Depending on the details of a particular model, flares may dominantly contribute to the diffuse neutrino flux (Murase et al., 2018). Besides, the atmospheric background is limited by the time window of a flare.
@@
Using publicly available data from SMARTS, Swift-UVOT, Swift-XRT, and Fermi-LAT, we obtain seven multi-wavelength quasi-simultaneous one-day averaged SEDs observed during the flare period. We describe the multi-wavelength data analysis in Sect. 2. We model (Sect. 3) the observable flare SEDs with the publicly-available AM3 program (Klinger et al., 2023, 2024) within three different frameworks: 1) assuming a standard leptonic model where the observed electromagnetic emission comes from relativistic electrons losing energy via synchrotron and inverse-Compton (IC) radiation in compact blobs passing through a quasi-stationary feature in the jet into which relativistic plasma is injected with large bulk Lorentz factor (“standing-feature” model, Sects. 3.1, 4.1); 2) utilizing the AM3’s capabilities of the time-dependent modelling, we attempt to model the whole flare in framework of a “moving-blob” model where the whole radiation comes from a single blob co-moving with the bulk motion of the relativistic jet plasma (“moving-blob” model, Sects. 3.2, 4.2); 3) assuming that in addition to electrons, protons might be co-accelerated in the jet with the same efficiency, we calculate the maximum-allowed by the data proton energy density $u^{\prime\mathrm{UL}}_{p}$ within a leptohadronic “standing-feature” model (Sects. 5, 6), providing our method to estimate the expected IceCube neutrino yield in Sect. 5.1 with corresponding results in Sect. 6.1. Our method to extrapolate the results for 3C 454.3 on the whole population of Fermi-LAT FSRQs is presented in Sect. 5.2 and the results of this are presented in Sect. 6.2. We discuss our findings in Sect. 7. In Sect. 8, we conclude that it is not possible to describe the seven days of the flare with a single “moving-blob” model, while a “standing-feature” model does it successfully. Moreover, we find that multi-wavelength flares of Fermi-LAT FSRQs similar to the studied 3C 454.3 flare contribute to the neutrino flux at energies $\geq 100$ TeV at the level of $\sim 0.5$% under the assumption that their emission can be described within a single-zone “standing-feature” leptohadronic model with the scaling between the observed integral neutrino and $\gamma$-ray fluxes inferred from the 3C 454.3 modelling for MJD 55517–55524.
@@ 2.1 Fermi-LAT – $\gamma$ rays
For our object of interest, 3C 454.3, we select the Fermi-LAT data in the energy range $100$ MeV $\leq E\leq$ 11 TeV and in time intervals corresponding to one day within the period MJD 55517–55524. The analysis is performed with the standard fermipy (Wood et al., 2017) package for the binned likelihood analysis. During the Fermi-LAT data analysis of the $15^{\circ}$ ROI (region of interest) centred at the position of 3C 454.3 we apply a cut on the zenith angle value $\theta_{z}\leq 90^{\circ}$ in order to remove contamination from the Earth’s limb. We use the P8R3_SOURCE_V3 instrument response function (evclass =128=128) with detected $\gamma$ rays both in FRONT and BACK of the Fermi-LAT (evtype =3=3) [^1] https://fermi.gsfc.nasa.gov/ssc/data/analysis/documentation/Cicerone/Cicerone_LAT_IRFs/IRF_overview.html. We divide the data into six bins per decade in energy, and the angular binning is $0.1^{\circ}$ per pixel. To describe the observed $\gamma$-ray emission we construct a model containing all sources from the 4th Fermi-LAT source catalogue 4FGL (Abdollahi et al., 2020) located within $15^{\circ}$ from the centre of the ROI2223C 454.3 is present itself in the 4FGL with the name 4FGL J2253.9+1609., the model of the diffuse galactic background gll_iem_v07, and the model of the isotropic $\gamma$-ray background iso_P8R3_SOURCE_V3_v1333https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html. We model the spectral shape of 3C 454.3 as a power law with a subexponential cutoff [^4] The Fermi-LAT Collaboration uses the term “superexponential cutoff” https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/source_models.html, but in our case b<1{b}<1, so we use the term “subexponential” instead. as it was modelled in the 4FGL catalogue:
$\frac{dN}{dE}=N_{0}\left(\frac{E}{E_{0}}\right)^{-\gamma_{1}}\exp{\left(-aE^{b}\right)}.$ (1)
The normalizations of the diffuse backgrounds, 3C 454.3 and other sources within $7.5^{\circ}$ from the centre of the ROI are left free during the initial fitting of the ROI with all spectral shapes fixed according to the catalogue; both normalizations and spectral shapes of other sources outside $7.5^{\circ}$ from the ROI centre are fixed to their catalogue values. After the initial fit, the spectral shape of 3C 454.3 is freed, and the fit is repeated. The results of the global Fermi-LAT SED fits are presented in Table 1. The $\gamma$-ray SED of 3C 454.3 is then obtained using the fermipy.GTAnalysis.sed method with the default parameter values [^5] https://fermipy.readthedocs.io/en/latest/advanced/sed.html#id3.
Parameter 55517 55518 55519 55520 55521 55522 55523 7-d-averaged
$N_{0}$ $3.22\pm 0.27$ $3.53\pm 0.07$ $3.60\pm 0.08$ $3.55\pm 0.08$ $1.67\pm 0.25$ $1.42\pm 0.05$ $1.22\pm 0.05$ $2.63\pm 0.26$
$\gamma_{1}$ $1.78\pm 0.05$ $1.952\pm 0.021$ $1.959\pm 0.022$ $1.806\pm 0.021$ $1.88\pm 0.08$ $1.96\pm 0.03$ $1.95\pm 0.04$ $1.889\pm 0.009$
$E_{0}$ $517.5$
$a$ $1.99$ $1.991$ $1.9908$ $1.9908$ $1.8$ $1.99076$ $1.9908$ $1.990755$
$\pm 0.27$ $\pm 0.003$ $\pm 0.0009$ $\pm 0.0018$ $\pm 0.5$ $\pm 0.00019$ $\pm 0.0016$ $\pm 0.000012$
$b$ $0.5184$
Table 1: Results of the Fermi-LAT $\gamma$-ray global SED fit in the energy range $100$ MeV $\leq E\leq$ 11 TeV obtained with fermipy for each of the seven days and the seven-day-averaged period. The columns indicate the start time of the one-day averaged observations in MJD. $N_{0}$ is expressed in $10^{-8}$ MeV-1 s-1 cm-2]; $E_{0}$ is in [MeV]; and $a$ is in $10^{-2}$ MeV-1].
Parameter 55517 55518 55519 55520 55521 55522 55523 7-d-averaged
obsID 00035030133 00035030134 00035030136 00035030138 00035030140 00035030142 00035030144 all shown obsIDs
00035030135 00035030137 00035030139 00035030141 00035030143 00035030145
$\kappa$ $5.3^{+4.2}_{-0.0}$ $5.3^{+4.2}_{-0.0}$ $5.3^{+2.9}_{-0.0}$ $5.3^{+3.0}_{-0.0}$ $9.5^{+0.0}_{-4.2}$ $9.5^{+0.0}_{-4.2}$ $8.3^{+1.2}_{-3.0}$ $7.7^{+1.8}_{-2.4}$
$N_{X}$ $19.4^{+5.1}_{-1.2}$ $3.3^{+2.9}_{-0.6} 22$$2.0^{+2.7}_{-0.5}$ $20.3^{+2.5}_{-0.5}$ $16.6^{+0.5}_{-1.9}$ $15.4^{+1.4}_{-1.8}$ $12.8^{+0.8}_{-1.3}$ $18.5^{+1.0}_{-1.2}$
$\Gamma_{1}$ $0.97^{+0.40}_{-0.28}$ $1.00^{+0.26}_{-0.09}$ $1.06^{+0.22}_{-0.07}$ $1.05^{+0.22}_{-0.08}$ $1.25^{+0.08}_{-0.35}$ $30^{+0.10}_{-0.42} 1.1$$0^{+0.15}_{-0.28} 1.1$$16^{+0.10}_{-0.16}$
$E_{\mathrm{br}}$ $1.6^{+0.9}_{-0.6}$ $1.74^{+0.44}_{-0.20}$ $2.09^{+0.79}_{-0.31}$ $2.0^{+0.5}_{-0.3}$ $2.0^{+0.7}_{-0.5}$ $1.6^{+1.5}_{-0.6}$ $1.9^{+0.5}_{-0.4}$ $1.96^{+0.25}_{-0.25}$
$\Gamma_{2}$ $1.69^{+0.25}_{-0.16}$ $1.58^{+0.08}_{-0.06}$ $1.60^{+0.11}_{-0.08}$ $1.54^{+0.10}_{-0.08}$ $1.69^{+0.15}_{-0.12}$ $1.56^{+0.16}_{-0.08}$ $1.64^{+0.14}_{-0.11}$ $1.61^{+0.04}_{-0.04}$
$\chi^{2}_{\mathrm{red}}$ $0.88$ $1.75$ $0.79$ $0.80$ $1.29$ $1.09$ $0.39$ $1.06$
Table 2: Results of the Swift-XRT X-ray SED fit obtained with ISIS for each of the seven days and the whole seven-day-averaged period. The asymmetric errors display the 6868% confidence interval estimated using conf_loop ISIS function. Parameter $\kappa$ for each single day is forced to be within 6868% confidence interval of the seven-day-averaged fit for the stability of the fits. The columns indicate the start time of the one-day-averaged observations in MJD. Parameter $\kappa$ is expressed in $10^{20}$ cm-2], $N_{X}$ is in $10^{-3}$ cm-2 s-1 keV-1], and $E_{\mathrm{br}}$ is in [keV].
Filter 55517 55518 55519 55520 55521 55522 55523 7-d-averaged
$v$ - - $12.7\pm 0.3$ - - - - -
$b$ - - $8.03\pm 0.19$ - - - - -
$u$ - - $5.10\pm 0.13$ $6.64\pm 0.14$ - - - -
$uw1$ - - $2.31\pm 0.05$ - - - $1.72\pm 0.04$ -
$um2$ - $1.60\pm 0.04$ $1.48\pm 0.04$ - - $1.241\pm 0.029$ - $1.44\pm 0.15$
$uw2$ $030\pm 0.023 - 1.1$$189\pm 0.029 -$ $0.940\pm 0.020$ - - $1.05\pm 0.10$
Table 3: Spectral flux densities and their uncertainties in mJy observed (before deredenning) with each filter by the Swift-UVOT for each of the seven days and the whole seven-day-averaged period. The columns indicate the start time of the one-day averaged observations in MJD. The Swift-UVOT uses the same obsIDs as the Swift-XRT (see Table 2). Filter $v$ corresponds to the wavelength of $5468$ Å, bb43924392 Å, uu34653465 Å, uw1uw126002600 Å, um2um222462246 Å, uw2uw219281928 Å.
@@ 2.2 Swift-XRT – X rays
The observed spectrum is modelled as a broken power law with the normalization $N_{X}$, break energy $E_{\mathrm{br}}$, and the two spectral indices ($\Gamma_{1}$ before the break and $\Gamma_{2}$ after the break) accounting for the photoelectric absorption along the LoS (line of sight) with the hydrogen column density $\kappa$. The details of the deabsorption procedure are presented in Appendix B. The results of the Swift-XRT SED fits are presented in Table 2.
@@ 2.3.1 Swift-UVOT
The Swift-UVOT SEDs are obtained with the swift-uvot-analysis-tools [^7] https://github.com/KarlenS/swift-uvot-analysis-tools program, which provides the values of the observed SED averaged over a given day within the time interval MJD 55517–55524. For the seven-day-averaged period, we calculate the arithmetic average optical/UV SED value in each filter, which was used at least three times within the period and incorporate the standard deviation of the flux values as an additional uncertainty, adding it quadratically to the statistical uncertainties (the same is done for the optical data from SMARTS and Steward Observatory).
@@ 2.3.3 Deredenning
The deredenning for IR, optical and UV data was performed using the extinction program (Barbary, 2016) assuming the extinction law of Fitzpatrick (1999), with the mean value of E(BV)E(B-V) extracted from IRSA (2022) using astroquery.ipac.irsa.irsa_dust module (Ginsburg et al., 2024, 2019) with the assumed value of $R_{V}=3.1$.